y + 9 = 1/4 (x -5)
y + 9 = 1/4x - 5/4
y + 1/4x - 10 1/4 (-41/4)
idk if this is right any help is appreciated thanks in advance

Answer by stanbon(75887)

y + 9 = 1/4 (x -5)
y + 9 = 1/4x - 5/4
------
y = (1/4)x - (5/4) - (36/4)
----
y = (1/4)x - (41/4)
=======================
Cheers,
Stan H.
========================